OpenStudy (anonymous):

1 MEDDAL WILL BE GIVEN PLEASE HELP!! Captain Jim is going to load his pirate ship with barrels. He has room for 12 barrels. He has allotted for 103 pounds of barrels on his ship. He can choose from 8-pound barrels (the variable a) or 11-pound barrels (the variable b). Which equations would be a constraint to Captain Jim maximizing his barrel load?

OpenStudy (anonymous):

ok i would set it up as \[a \times 12\le103\]\[b \times12\le103\]

OpenStudy (anonymous):

but it gives me this options...a. 8a = 103 - 11b b. 8a = 103 + 11b c. a - b = 12 d. a = 12 + b

OpenStudy (anonymous):

ok i would say that either d or a. but i am not sure. @Salmon which do you think it is

OpenStudy (anonymous):

8a+11b=< 103 a+b=<12

OpenStudy (anonymous):

so it's c? a. 8a = 103 - 11b b. 8a = 103 + 11b c. a - b = 12 d. a = 12 + b

OpenStudy (anonymous):

Its a cheak option d...it would be a=12-b because option c and d are same...

OpenStudy (anonymous):

so its D?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

A